OpenStudy (anonymous):

Let h and g be two functions defined by h(x)=4x^2-5 g(x)=7x-2 for any real number x . Find h/g(-4) and then find all values that are NOT in the domain of h/g .

OpenStudy (anonymous):

h(x)/g(x) =(4x^2-5)/(7x-2) Now, h/g(-4) =(4(-4)^2 -5) / (7(-4) -2)

OpenStudy (anonymous):

Simplify it.

OpenStudy (anonymous):

59/26

OpenStudy (anonymous):

?

OpenStudy (anonymous):

59/(-30)

OpenStudy (anonymous):

damn haha okay now what

OpenStudy (anonymous):

Now, h/g is undefined when denominator is 0 SO, find for what value of x the denominator will be 0

OpenStudy (anonymous):

@kaelyb can u?

OpenStudy (anonymous):

hmm i dont know what it would be!

OpenStudy (anonymous):

for what value of x 7x-2 is 0

OpenStudy (anonymous):

Hint: 7x-2=0 Now, solve for x

OpenStudy (anonymous):

ohhh gotchu haha 2/7

OpenStudy (anonymous):

So, when x=2/7 the denominator is 0. so 2/7 is not in the domain of h/g

OpenStudy (anonymous):

okay cool thanks!

OpenStudy (anonymous):

welcome